OpenStudy (anonymous):

Using complete sentences, explain the difference between a linear equation and a quadratic equation.

OpenStudy (anonymous):

Linear Function: A function with no exponents other than one and with no products of the variables (eg, y=x+4, y= -4, and 3x-4y = 1/2 are linear functions); in a rectangular coordinate system, the graph of a linear function is a line. It is a function that has a constant rate of change. It has a graph which is a straight line.Such a function can be written as f(x) = mx + b, where m and b are real constants and x is a real variable. Quadratic function: A quadratic function, in mathematics, is a polynomial function of the form y = ax^2+bx+c,where, a is not equal to "0",where a, b, and c are arbitrary, but fixed. A quadratic function is also referred to as a degree 2 polynomial or a 2nd degree polynomial, because the highest exponent of x is 2. The graph of this function is a parabola.

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

no problem, medal thanks :D

OpenStudy (anonymous):

How can I word this different and shorters than that

jimthompson5910 (jim_thompson5910):

Linear equation: graphs a straight line Quadratic equation: graphs a "bowl" shaped curve That's the basic visual difference

jimthompson5910 (jim_thompson5910):

In terms of algebra, a linear equation has one x-intercept and this is always true. On the other hand, a quadratic equation can have 0, 1 or 2 x-intercepts (depending on the equation)

OpenStudy (anonymous):

What do the equation look like?

jimthompson5910 (jim_thompson5910):

Linear equation: 

OpenStudy (anonymous):

written

jimthompson5910 (jim_thompson5910):

Ex: Linear equation: y = 2x+3 Quadratic: y = x^2+4x-5

OpenStudy (anonymous):

Whats the Quadrtic with the letters?

OpenStudy (anonymous):

\[Quadratic : y = x ^{2} + 4x -5\]

jimthompson5910 (jim_thompson5910):

oh it's y = ax^2+bx+c

OpenStudy (anonymous):

y = ax^2+bx+c is the base equation.

OpenStudy (anonymous):

just like the linear equation y = ax + b